In a scene from the movie, Superman was shown running after a train. The train moves at 54 m/s. How much work must he do in order to stop the train? Mass of train = 425 000 kg.



1
Expert's answer
2021-02-24T12:45:44-0500
"\Delta KE=W,""KE_f-KE_i=W,""0-\dfrac{1}{2}mv^2=W,""W=-\dfrac{1}{2}\cdot425000\ kg\cdot(54\ \dfrac{m}{s})^2=-6.19\cdot10^8\ J."

The Superman must do "6.19\cdot10^8\ J" of work in order to stop the train. The sign minus means that Superman pushes the train in the direction opposite to the motion of the train.